OpenStudy (anonymous):

What is the total cost if you buy 3 goldfish for $1.90 each, 3 angelfish for $6.10 each, and 12 neon tetras for $1.53 each?

OpenStudy (anonymous):

1.90*3=5.70.6.10*3=18.30.1.53*12=18.36. Add those all together and your answer should be 42.36.

OpenStudy (anonymous):

1.90 * 3 6.10 * 3 1.53 * 12

OpenStudy (anonymous):

find the product of each one and add them up.

OpenStudy (anonymous):

1.90 * 3 = 5.7 6.10 * 3 = 18.3 1.53 * 12 = 18.36

OpenStudy (anonymous):

5.7 + 18.3 + 18.36 = 42.36

OpenStudy (anonymous):

42.36

OpenStudy (anonymous):

Precisely :) Good Job :D

OpenStudy (anonymous):

THANKS

OpenStudy (anonymous):

yw :)